OpenStudy (anonymous):

how do i write an equation of the line parallel to the line y= 2x+5 and passing through the point (4,6)?

OpenStudy (anonymous):

what do you know about the relationship between parallel lines?

OpenStudy (anonymous):

parallel = same slope y= 2x + b 6 = 2(4) + b 6 = 8 + b b = -2 eq: y= 2x -2

OpenStudy (anonymous):

particularly their slopes?

OpenStudy (anonymous):

or you can just have some random scrandom just give you the answer thanks bud

OpenStudy (anonymous):

ok ok how can b=-2 if b is said to equal 5?

OpenStudy (anonymous):

because the b value of parallel lines are not the same. If you think of parallel lines as just the same line in a different position on the graph then the change in b values is what causes the change in position otherwise they are the exact same line.

OpenStudy (anonymous):

ok thank you now i understand I just started this business algebra and I haven't done this type of math in over 7 years.

OpenStudy (anonymous):

Its fine your welcome